OpenStudy (jenniferjuice):

simpson drives his car with an average velocity of 48km/h to the east.How long will it take him to drive 144 km on a straight highway?

OpenStudy (b2st):

Okay let's start figuring out what we have.

OpenStudy (b2st):

So what is 48km/h? & 144km ? What goes where? x= ? t=? avg= ?

OpenStudy (b2st):

>_> I see you're not replying eh? welp, you should use the Average Velocity formula and find what you are missing! Plug in the numbers you have and find what you need.

OpenStudy (jenniferjuice):

sorry mom made me throw out the trash! but im back!!!

OpenStudy (b2st):

Oh okay! Let's do this problem.

OpenStudy (b2st):

So tell me what goes where?

OpenStudy (jenniferjuice):

well x=48

OpenStudy (b2st):

No that's incorrect.

OpenStudy (jenniferjuice):

avg=48

OpenStudy (b2st):

x is the displacement btw.

OpenStudy (b2st):

Yes, that's correct.

OpenStudy (jenniferjuice):

x=144

OpenStudy (b2st):

Right! So what formula do you think we should use?

OpenStudy (b2st):

(unless you physics teacher didn't teach you about this )

OpenStudy (jenniferjuice):

Varg*Δt=Δx

OpenStudy (b2st):

well. The original problem for average velocity would be Avg = Δx / Δt

OpenStudy (b2st):

^ keep that in mind.

OpenStudy (b2st):

So what are looking for?

OpenStudy (jenniferjuice):

okay yeah i got that equation from there

OpenStudy (b2st):

From the previous, what are we missing?? We got x & the avg. What do we need?

OpenStudy (jenniferjuice):

time!

OpenStudy (b2st):

Yes! :D

OpenStudy (b2st):

So let's rearrange Avg = Δx / Δt

OpenStudy (b2st):

Try to make Δt alone.

OpenStudy (jenniferjuice):

Δx*Avg=Δt ?

OpenStudy (b2st):

Incorrect. Look carefully.

OpenStudy (b2st):

You can't multiply Δx, because Δx is in the numerator.

OpenStudy (b2st):

Although you were correct on multiplying, but the wrong part.

OpenStudy (jenniferjuice):

so... Δx/Avg ?

OpenStudy (b2st):

Yes correct!

OpenStudy (b2st):

so you have Δt = Δx / Avg

OpenStudy (b2st):

All you do is plugin and you get your answer. I hope this help! :]

OpenStudy (jenniferjuice):

is it 3?

OpenStudy (b2st):

Yes ! 3 hours. :]

OpenStudy (jenniferjuice):

okay thanks! i haven't actually understood this till now! its actually really easy ! :D